# Identifying Compton-thick AGNs in the COSMOS

### v1

@@ 6 Summary
In this work, we aim to identify CT-AGNs hidden in AGNs whose X-ray flux is below the current flux limits in the COSMOS. First, we selected a sample of 1,104 MIR-selected AGNs that were covered but individually undetected by X-ray. Then, we reduced the Chandra observational data covering the field and derived upper limits on the X-ray luminosity for all AGNs in our sample. We subsequently collected multiwavelength photometric data for the sample and extracted key physical parameters for each source through SED fitting, such as 6 m luminosities of AGNs, stellar masses, and SFRs. After that, through MIR diagnosis, 7 to 23 sources could be identified as CT-AGNs. To confirm that our CT-AGNs are indeed heavily absorbed in the X-ray band, we performed a stacking analysis. The stacked signal is detected at 3 significance in the soft band, with a flux of , whereas the hard-band excess is only 1 significant, corresponding to . Fitting the stacked soft- and hard-band fluxes with the zphabszpowerlw+constantzpowerlw+pexmon model obtained two parameter regions that are consistent with the data at the 95% confidence level , corresponding to Compton-thin and Compton-thick solutions. The Compton-thin solution can be ruled out, leaving the Compton-thick absorber as the only viable scenario and confirming these sources as CT-AGNs. We fitted their soft- and hard-band fluxes with the zphabszpowerlw+constantzpowerlw+pexmon model, and found that when Compton-thick absorption was included did the model provide a satisfactory fit to both bands. This provides direct and compelling evidence that these sources are CT-AGNs. Although we have already identified some CT-AGNs in the MIR-selected AGNs, a considerable population of CT-AGNs was still missed by our selection. Therefore, we estimated and discussed the number of missed CT-AGNs in our sample and proposed a practical strategy to search for them. Finally, comparing host-galaxy properties between CT-AGNs and non-CT-AGNs, we did not find evidence for a correlation between intense star formation and CT-AGNs.

### v2

@@ 6 Summary
In this work, we aim to identify CT-AGNs hidden in AGNs whose X-ray flux is below the current flux limits in the COSMOS. First, we selected a sample of 1,104 MIR-selected AGNs that were covered, but individually undetected in X-ray. Then, we reduced the Chandra observational data covering the field and derived upper limits on the X-ray luminosity for all AGNs in our sample. We subsequently collected multiwavelength photometric data for the sample and extracted key physical parameters for each source through SED fitting, such as six m luminosities for AGNs, stellar masses, and SFRs. In the next step, using an MIR diagnosis, 7 to 23 sources could be identified as CT-AGNs. To confirm that our CT-AGNs are indeed heavily absorbed in the X-ray band, we performed a stacking analysis. The stacked signal was detected at 3 significance in the soft band, with a flux of , whereas the hard-band excess is only 1 significant, corresponding to . Fitting the stacked soft- and hard-band fluxes with the zphabszpowerlw+constantzpowerlw+pexmon model allowed us to obtain two parameter regions that are consistent with the data at the 95% confidence level, corresponding to Compton-thin and Compton-thick solutions, respectively. The Compton-thin solution can be ruled out, leaving the Compton-thick absorber as the only viable scenario and confirming these sources as CT-AGNs. We fit their soft- and hard-band fluxes with the zphabszpowerlw+constantzpowerlw+pexmon model and found that when Compton-thick absorption was included, the model did provide a satisfactory fit to both bands. This provides direct and compelling evidence that these sources are CT-AGNs. Although we have already identified some CT-AGNs in the MIR-selected AGNs, a considerable population of CT-AGNs was still missed by our selection. Therefore, we have estimated and discussed the number of missed CT-AGNs in our sample and proposed a practical strategy to search for them. Finally, based on our comparison of host-galaxy properties between CT-AGNs and non-CT-AGNs, we did not find any evidence to support a correlation between intense star formation and CT-AGNs.